# Efficient optimization of neural network backflow for ab-initio quantum chemistry

### v1

@@ III.1 Benchmarks
Figure 3: Experiments comparing the impact of batch size $N_{s}$ on energy improvements for Li2O, using the STO-3G basis set with canonical HF orbitals. Results are shown for the original algorithm from our previous work [34] and our current improved algorithm. The energy error decreases approximately following $O(N_{s}^{-1.979616\pm 0.006350})$ with $R^{2}=0.935316$ for the improved algorithm introduced in this work and $O(N_{s}^{-1.166309\pm 0.014370)})$ with $R^{2}=0.880579$ for the algorithm used in Ref. 34. Here the linear regression is done by the weighted least square method since the reported energies have uncertainties.
@@ III.2 Ablation study
Here we demonstrate how each algorithm introduced in Sections II.2, II.3, II.4, and II.5 improves the energy accuracy and computational efficiency through an ablation study on Li2O with the STO-3G basis set and HF orbitals. We cumulatively add these enhancements to our previous algorithm [34] to highlight their individual and combined contributions (see Figure 4).
Figure 4 shows that the introduction of the Gumbel top-$k$ trick alone reduces the energy error by two orders of magnitude without additional computational cost. In contrast to Ref. 34, which uses the union of the core space $\mathcal{V}$ and the entire connected space $\mathcal{C}$ as the total space $\mathcal{U}$, our preselection step generates a more compact important subspace $\tilde{\mathcal{U}}$ that is $\frac{1}{N_{o}-N_{e}}$ times smaller than $\mathcal{U}$. This reduces the number of amplitude evaluations by the same factor, decreasing the per-optimization-step time from 0.47 to 0.28 seconds.
@@
The truncated local energy evaluation strategy further accelerates training by replacing full neural network evaluations with simpler $O(\log{\absolutevalue{\tilde{\mathcal{U}}}})$ ) lookups, improving computational efficiency (the per-iteration time is reduced from 0.28 to 0.07 seconds) without compromising energy accuracy.
Finally, incorporating physical knowledge into the model architecture enhances the expressiveness of the NNBF ansatz and further improves energy accuracy, all while maintaining the same computational complexity. Although integrating spin-flip symmetry increases the wall time per optimization step by a factor of 1.42, the scaling remains unchanged.
@@
Altogether, these enhancements enable us to achieve significantly better energy accuracy and considerably reduce computational time for a given batch size and network architecture.
@@
Figure 4: Evaluation of accuracy and speed improvements achieved through the algorithmic enhancements detailed in Section II.4, II.2, II.3, and II.5 for training Li2O with the STO-3G basis set and canonical HF orbitals. The experiments use a fixed batch size of $N_{s}=1024$. Each bar represents the average energy error and per-iteration wall time across different seeds, with individual points corresponding to separate runs of the same model and method. The algorithmic improvements are applied cumulatively
@@ III.3 NNBF Expressiveness and the Role of IPR
Figure 5: Investigating the relationship between NNBF expressiveness and the inverse participation ratio (IPR) of the molecular ground state through the dissociation curves of (a) CH4 and (b) N2 molecules, using the STO-3G basis set with canonical HF orbitals. For each set of molecule geometry and network architecture, three trials are performed with different random seeds; each data point shows the smallest absolute relative energy error among them. The relative energy error is computed as the difference between the total NNBF energy and the total FCI energy, divided by the electronic FCI energy (excluding the fixed, non-variational nuclear repulsion contribution). The absolute relative energy error (left y-axis) and ground-state IPR (right y-axis) are plotted as functions of interatomic distance. To eliminate approximation errors in energy and its gradient, optimization is performed over the entire Hilbert space ($N_{s}=N_{t}$). Double precision is employed for exact inference energy calculations, as higher precision becomes crucial when the NNBF state closely approximates the true ground state. Different hidden unit sizes are used to illustrate overparameterized ($h=64$) and underparameterized ($h=16,32$) scenarios.
@@ IV Conclusions
In this work, we have demonstrated that the proposed algorithmic enhancements—preselection, truncated local energy evaluation, Gumbel top-$k$ sampling, and physics-informed encoding—significantly improve the performance of the NNBF ansatz across various molecular systems and strongly correlated problems. Our method consistently achieves lower ground-state energies than existing NQS approaches and surpasses conventional CCSD calculations, particularly for larger molecules where traditional methods struggle. When FCI data is available, NNBF often attains results closer to the exact solution than CCSD(T). In the challenging frozen-core N2 system with the cc-pVDZ basis, NNBF achieves variational energies competitive with or better than state-of-the-art SCI and FCIQMC methods. The ablation study highlights how each enhancement contributes to both improved energy accuracy and reduced computational cost, with the Gumbel top-$k$ trick alone reducing the energy error by two orders of magnitude without additional overhead and the combination of preselection and truncated local energy strategies further reducing the per-iteration time significantly.
A direct comparison with our previous method [34] further emphasizes the efficiency gains achieved through these improvements. While both approaches exhibit polynomial scaling of energy accuracy with batch size, the updated algorithm achieves a notably faster convergence rate and requires orders of magnitude fewer samples to reach the same level of precision. For example, in the Li2O benchmark, the improved method attains similar accuracy with a batch size hundreds of times smaller than that required by the previous approach. This reduction in sample complexity, combined with decreased wall time per optimization step, underscores the effectiveness of the proposed techniques in enhancing computational efficiency. Lastly, our analysis of NNBF expressiveness shows that lower IPR values generally correspond to higher relative energy errors, indicating that more delocalized states are harder to approximate. Despite this trend, factors beyond IPR also influence the optimization difficulty, highlighting the complexity of the amplitude landscape.
@@ Appendix A Experimental Setup
The internal neural network used in this work is a multilayer perceptron (MLP) with $h$ hidden units, $L$ hidden layers, and $D$ backflow determinants, where residual connections are incorporated when $L>1$. We use the AdamW optimizer to minimize the energy expectation value of our NNBF ansatz and approximate the ground state of various molecules. The default hyperparameters are listed in Table 3. Baseline energies for HF, CCSD, CCSD(T), and FCI calculations are obtained using the PySCF software package [48]. Each calculation was performed using 1 NVDIA V100 gpu.
After training with the algorithmic enhancements described in Sections II.2, II.3, II.4, and II.5, we perform a separate MCMC inference step to estimate the variational energy of the trained NNBF state. This inference procedure follows the approach detailed in Appendix A of Ref. 34, providing an unbiased stochastic estimate of the energy. These values are reported in all experiments, except when investigating the relationship between NNBF expressiveness and the IPR, where the energy is computed exactly over the full Hilbert space.
Hyperparameter Value
Energy unit Hartree
Framework JAX
Precision float32
Optimizer AdamW
AdamW’s $\beta_{1}$ 0.9
AdamW’s $\beta_{2}$ 0.999
AdamW’s $\epsilon$ $1\times 10^{-8}$
AdamW’s $\lambda$ 0.0001
Learning rate $10^{-3}\times(1+10^{-4}t)^{-1}$
Training iterations 100000
Number of pretraining walkers 8192
Pretraining iterations 500
Number of dominant configurations 8
for walker initialization
MCMC burn-in steps $20\times 2N_{e}$
right after initialization
Number of MCMC walkers 1024
MCMC downsample interval $1\times 2N_{e}$
MCMC iterations 1000
Number of hidden layers 2
Hidden units 256
Number of determinants 1
Preselection reduction factor $1/(N_{o}-N_{e})$
@@
Table 3: Default hyperparameters used for all experiments in the paper, unless explicitly stated otherwise
symbol description
$N_{e}$ number of electrons
$N_{o}$ number of spin-orbitals
$N_{t}=\binom{N_{o}}{N_{e}}^{2}$ number of physically valid configurations
$N_{s}$ number of samples
$L$ number of hidden layers
$h$ hidden units
$D$ number of determinants
$N_{w}$ number of MCMC walkers
Table 4: Table of notations

### v2

@@ III.1.3 All-electron and frozen-core N2 calculations
Method (Parameter) 2.118 Bohr 2.7 Bohr 3.6 Bohr
RHF -108.949378 -108.833687 -108.767549
CCSD -109.267626 -109.131491 -108.975885
CCSD(T) -109.28030 -109.150645 -108.982836
CCSDT -109.280323 -109.156703 -108.990518
CCSDTQ -109.281943 -109.162264 -108.993736
DMRG (m=1000) -109.281878 -109.163087 -108.997549
DMRG (m=2000) -109.282088 -109.163467 -108.997939
DMRG (m=4000) -109.282157 -109.163572 -108.998052
NNBF (|$\absolutevalue{\mathcal{S}}$ |=65536 |$\absolutevalue{\mathcal{V}}$ |=16384) -109.282036(48) -109.163438(67) -108.997789(142)
Table 3: Ground-state energy benchmark for the all-electron N2 molecule (cc-pVDZ basis set) at three bond lengths: 2.118, 2.7, and 3.6 Bohr. Results from our NNBF method are compared against conventional quantum chemistry methods (HF, CCSD, CCSD(T), CCSDT, and CCSDTQ) and state-of-the-art Density Matrix Renormalization Group (DMRG) calculations [39]. The DMRG calculations were performed using canonical UHF orbitals with several bond dimensions ($m=1000,2000,\text{and }4000$). The NNBF calculations, in contrast, used CCSD natural orbitals. Each reported NNBF energy is variational, obtained from a single post-training MCMC inference with a $(D,L,h)=(1,2,512)$ network architecture.
@@ III.2.1 Cumulative Feature Addition
This subsection demonstrates how each algorithmic enhancement from Sections II.2, II.3, II.4, and II.5 contributes to improvements in energy accuracy and computational efficiency. The analysis is performed via an ablation study on the Li2O molecule (STO-3G basis, starting from HF orbitals). These enhancements are cumulatively added to our previous algorithm [34] to highlight their individual and combined contributions, with key results illustrated in Figure 4.
The most significant improvement in accuracy comes from introducing the Gumbel top-$k$ trick (Section II.4). As shown in Figure 4, this sampling method alone reduces the energy error by two orders of magnitude without imposing additional computational cost per step.
Efficiency is first enhanced by the Intermittent Target Selection (ITS) method (Section II.2). In contrast to our previous work [34], which used the entire connected space $\mathcal{C}$ as the target space, ITS constructs a much more compact subspace $\mathcal{U}$. For the settings studied, its size is reduced by a factor of $l\sim N_{o}-N_{e}$. This directly lowers the number of required amplitude evaluations, reducing the time per optimization step from 0.327 to 0.185 seconds for the system tested.
@@
The training process is further accelerated by our truncated local energy strategy (Section II.3). By replacing many computationally demanding neural network evaluations with efficient $O(\log{|\mathcal{U}|})$ ) lookups of precomputed amplitudes, this strategy significantly boosts computational efficiency. In our example, this modification reduced the per-optimization time from 0.185 to 0.060 seconds, without compromising energy accuracy.
Finally, incorporating physical knowledge into the model architecture (Section II.5) enhances the expressiveness of the NNBF ansatz, leading to another gain in energy accuracy. These features maintain the same asymptotic computational complexity. However, we note that enforcing spin-flip symmetry increases the wall time by a constant factor (approx. 1.58x in our tests), leaving the overall scaling unchanged. While the orbital envelope offers a marginal improvement for this specific system, we have found it is crucial for achieving high accuracy in larger molecules.
@@
Collectively, these algorithmic enhancements enable the achievement of significantly improved energy accuracy and considerably reduced computational time for a given batch size and network architecture.
@@
Figure 4: Evaluation of accuracy and speed improvements achieved through the algorithmic enhancements detailed in Section II.4, II.2, II.3, and II.5. Each bar represents the average energy error and per-optimization-step wall time over three independent runs, with the individual data points also shown. All calculations were performed on the Li2O molecule with the STO-3G basis set, starting from canonical HF orbitals. The experiments used a fixed batch size of $|\mathcal{S}|=|\mathcal{V}|=1024$ and a network architecture of $(D,L,h)=(1,2,64)$. The algorithmic improvements are applied cumulatively.
III.2.2 The effect of Gumbel noise, inclusion probability, and renormalization
Given that Gumbel top-$k$ selection can significantly improve accuracy (as shown in Figure 4), it is instructive to investigate the contribution of its individual components. To isolate these contributions, we performed an ablation study on the Li2O molecule (STO-3G basis, $|\mathcal{S}|=|\mathcal{V}|=1024$). We systematically enabled or disabled three key features of the sampling and weighting scheme: (1) the use of Gumbel noise to perturb selection probabilities, (2) the use of the inclusion probability $q_{i}(\kappa)$ [Eq. (10)] in the assignment of importance weights, and (3) the subsequent renormalization of these weights.
The findings, presented in Figure 5, demonstrate the distinct role of each component. Adding only Gumbel noise to the sampling process, without the corresponding unbiased reweighting, provides a marginal but noticeable advantage. This is expected, as the estimators for the energy and its gradient [Eqs. (1) and (3)] still depend entirely on the specific configurations in the sample $\mathcal{S}$. Conversely, incorporating the inclusion probability $q_{i}(\kappa)$ into the importance weights is crucial, reducing the energy error by an order of magnitude even without renormalization. Applying renormalization to these weights provides a final, noticeable improvement to the energy.
These observations align with the description in Section II.4. The inclusion probability provides an unbiased estimator over the entire target space $\mathcal{U}$, while the subsequent renormalization, in practice, improves the estimation by reducing variance, albeit at the potential cost of introducing a small bias.
Figure 5: Ablation study showing the individual contribution of each component of the Gumbel top-$k$ sampling method (introduced in Section II.4) to the overall energy improvement detailed in Figure 4. All calculations were performed on the Li2O molecule (STO-3G basis) with a $(D,L,h)=(1,2,64)$ network architecture and parameters $|\mathcal{S}|=|\mathcal{V}|=1024$. The labels on the x-axis indicate the components used in each experiment: Gb refers to the use of Gumbel noise during sample selection; Ip signifies that the importance weights incorporate the inclusion probability ($q_{i}(\kappa)$); and Nm denotes that the final weights are renormalized.
III.2.3 Local energy approximation strategy comparison
In Section II.3, we claimed that our truncated local energy strategy yields a more accurate approximation, and thus better training performance, than commonly used approaches at a comparable cost. Our method, denoted $E_{l}(\mathbf{x}_{i}|\mathcal{U})$, calculates the local energy for each sample $\mathbf{x}_{i}\in\mathcal{S}$ by summing over connected configurations within the entire target space $\mathcal{U}$. This contrasts with the common strategy, here denoted $E_{l}(\mathbf{x}_{i}|\mathcal{S})$, which restricts this sum to the much smaller sample set $\mathcal{S}$ itself [32, 35, 36]. The common approach typically also defines the importance weights as the amplitude-squared values renormalized within $\mathcal{S}$ (i.e., $w_{i}=|\psi(\mathbf{x}_{i})|^{2}/\sum_{\mathbf{x}_{j}\in\mathcal{S}}{|\psi(\mathbf{x}_{j})|^{2}}$). This combination of local energy truncation and weighting makes the training objective variational with respect to the sample set $\mathcal{S}$.
To substantiate our claim and isolate the sources of improvement, we performed a comparative study on the Li2O molecule (STO-3G basis) across various batch sizes ($|\mathcal{S}|=|\mathcal{V}|$). We designed a cumulative comparison starting from a baseline that represents the common approach (Method 0): using $E_{l}(\mathbf{x}_{i}|\mathcal{S})$ with simple amplitude-squared weights and reporting the best objective function value from training. We then assess the impact of three modifications in sequence. First (Method 1), we leave the wave-function generated from Method 0 but report the final energy using a global post-training MCMC inference which estimates the energy of the full wave-function. Second (Method 2), we additionally improve the training by incorporating the inclusion probability, $q_{i}(\kappa)$ [Eq. (10)], into the importance weights. Finally (Method 3), our full approach combines these improvements with our more accurate local energy calculation, $E_{l}(\mathbf{x}_{i}|\mathcal{U})$. The results of this study are depicted in Figure 6.
The results in Figure 6 reveal three key insights. First, comparing Method 1 to Method 0 shows that simply reporting the post-training MCMC inference energy provides a more accurate global estimate than using the best training objective value, reducing the error by an average factor of 2.12. This implies that the NNBF learns information about the wavefunction beyond the subspace it is directly trained on, reaffirming the representability and learnability of the ansatz.
Second, comparing Method 2 to Method 1 demonstrates that introducing the inclusion probability to the importance weights significantly improves performance, reducing the energy error by a further factor of 3.23. This confirms our conclusion from Section III.2.2 that proper reweighting is crucial for improving the objective function estimation when using stochastic sampling without replacement.
Third, comparing Method 3 (our full approach) to Method 2 shows that switching the local energy calculation from the sample set ($E_{l}(\mathbf{x}_{i}|\mathcal{S})$) to the target space ($E_{l}(\mathbf{x}_{i}|\mathcal{U})$) yields another non-trivial improvement, reducing the remaining error by an average factor of 2.94. This substantiates our central claim that leveraging the larger and more significant subspace $\mathcal{U}$ provides a more accurate local energy approximation, leading to superior overall performance. Importantly, we verified that the computational cost per optimization step is comparable for each method, confirming that these improvements in accuracy are not achieved at the expense of increased per-step complexity.
Figure 6: Cumulative performance comparison of different energy estimation strategies. All calculations were performed on the Li2O molecule using the STO-3G basis set with canonical Hartree-Fock (HF) orbitals and a $(D,L,h)=(1,2,64)$ network architecture. The data points for each batch size correspond to three independent runs, and the accompanying lines are least-squares fits to this data. The four methods compared add features cumulatively. Method 0 (baseline) uses a local energy sum over the sample set, $E_{L}(\mathbf{x}_{i}|\mathcal{S})$, with simple amplitude-squared weights, and reports the best training objective energy. Method 1 uses the same training but reports a post-training MCMC inference energy. Method 2 further improves the weights by incorporating the inclusion probability ($q_{i}(\kappa)$). Method 3 (our full approach) combines these improved weights with our more accurate local energy calculation, $E_{L}(\mathbf{x}_{i}|\mathcal{U})$. The energy error of Method 3 decreases with the sample size $N=|\mathcal{S}|$ approximately following $N^{-1.038(69)}$ ($R^{2}=0.859$).
@@ III.3 NNBF Expressiveness and the Role of IPR
[…]
@@ IV Conclusions
In this work, we have demonstrated that our proposed algorithmic enhancements—Intermittent Target Selection (ITS), truncated local energy evaluation, Gumbel top-$k$ selection, and physics-informed encoding—significantly improve the performance of the NNBF ansatz. Across a range of molecular systems, our method consistently achieves lower ground-state energies than existing NQS approaches and surpasses conventional CCSD and CCSD(T) calculations. The NNBF demonstrates its ability to capture both dynamic and static correlation by accurately matching the full dissociation curve of H2O in both near-equilibrium and highly-stretched regimes. Furthermore, for the challenging N2 molecule, our method achieves variational energies competitive with those from state-of-the-art SCI, FCIQMC, and DMRG calculations.
Our ablation studies quantify the impact of each new technique. We show that Gumbel top-$k$ sampling alone reduces the energy error by two orders of magnitude with no computational overhead, while the combination of ITS and our truncated local energy strategy significantly reduces the per-iteration wall time. We also present two additional focused comparative studies to provide deeper insight. The first dissects the Gumbel top-$k$ method, isolating the contributions from the stochastic noise, the inclusion probability weighting, and the final weight renormalization. The second study systematically compares our local energy strategy against the commonly used alternative, substantiating the benefits of our approach.
Lastly, our analysis of NNBF expressiveness shows that lower IPR values generally correspond to higher relative energy errors, indicating that more delocalized states are harder to approximate. Despite this trend, factors beyond IPR also influence the optimization difficulty, highlighting the complexity of the amplitude landscape.
@@ A.1 Training and Energy Inference Procedure
The internal neural network used in this work is a multilayer perceptron (MLP) with $h$ hidden units, $L$ hidden layers, and $D$ backflow determinants, where residual connections are incorporated when $L>1$. We use the AdamW optimizer to minimize the energy expectation value of our NNBF ansatz and approximate the ground state of various molecules. The default hyperparameters are listed in Table 4. Baseline energies for HF, CCSD, CCSD(T), and FCI calculations are obtained using the PySCF software package [49].
After training with the algorithmic enhancements described in Sections II.2, II.3, II.4, and II.5, some of which may be turned off in ablation studies, the energy expectation value and its statistical uncertainty are determined through post-training inference using a Markov Chain Monte Carlo (MCMC) procedure. We employ $N_{w}=1024$ concurrent walkers, each generating a Markov chain by sampling from the unnormalized probability distribution $\bar{p}_{\theta}(\bm{x})=\psi_{\theta}(\bm{x})^{2}$ using the Metropolis–Hastings algorithm. Proposed moves consist of swapping an occupied spin-orbital with an unoccupied one of the same spin, and to reduce autocorrelation, the chains are downsampled at an interval of $K_{1}=10N_{e}$ iterations.
Before this sampling begins, the walkers are initialized to ensure robust exploration and avoid trapping in local minima, following the ensemble method outlined in ref. 50. Specifically, initial positions are drawn from a distribution defined by the eight most dominant configurations identified in the final stage of training. The walkers then undergo a burn-in period of $K_{2}=100K_{1}$ iterations to reach equilibrium.
Following the burn-in, we collect $T=1000$ configurations from each walker. The final energy is computed as the mean over the total $T\times N_{w}$ collected samples, and the reported uncertainty is the standard error of this mean, given by $\sqrt{\text{Var}(E)/(T\times N_{w})}$. The energies reported in all experiments are obtained via this MCMC procedure, except when investigating the relationship between NNBF expressiveness and the IPR, where the energy is computed exactly over the full Hilbert space.
Parameter / Symbol Description Default Value / Definition
A. Physical System & Configuration Spaces
$N_{e}$ Number of electrons Problem-dependent
$N_{o}$ Number of spin-orbitals Problem-dependent
$\mathcal{H}$ The physical Hilbert space Problem-dependent
$\mathcal{V}$ Core space See Section II.2
$\mathcal{C}$ Connected space Generated from core space
$\mathcal{U}$ Target space See Section II.2
$\mathcal{S}$ Sample set for energy estimations See Section II.4
B. Model Architecture (MLP)
$D$ Number of backflow determinants 1
$L$ Number of hidden layers 2
$h$ Number of hidden units per layer 256
C. Optimizer & Training Schedule
Optimizer Algorithm for parameter updates AdamW
$\beta_{1},\beta_{2}$ AdamW exponential decay rates 0.9, 0.999
$\epsilon$ AdamW epsilon for numerical stability $1\times 10^{-8}$
$\lambda$ AdamW weight decay $1\times 10^{-4}$
Learning Rate ($t$) Initial rate with decay over iteration $t$ $10^{-3}\times(1+10^{-5}t)^{-1}$
Pretraining Iterations Number of steps before main training 500
$N_{w}$ (pretraining) Number of walkers during pretraining 8192
D. Algorithmic Enhancements
$l$ Speedup factor for ITS $N_{o}-N_{e}$
E. Implementation Details
Framework Core computational library JAX
Precision Floating-point precision float32
Energy Unit Standard unit for energy values Hartree
F. Post-Training MCMC Inference
$M_{\text{init}}$ Dominant configurations for initialization 8
$N_{w}$ (inference) Number of MCMC walkers 1024
$K_{2}$ Burn-in steps per walker $100K_{1}$
$K_{1}$ Downsample interval (iterations) $10N_{e}$
$M$ Samples collected per walker 1000
@@ F. Post-Training MCMC Inference
Table 4: Consolidated hyperparameters and notations used for all experiments, unless explicitly stated otherwise.
A.2 Geometry and hyperparameters used for Section III.1.1
This section provides the specific molecular geometries and key training hyperparameters used for the ground-state energy calculations presented in Section III.1.1. The following table lists the Cartesian coordinates (in Angstroms) for each molecule. Alongside each geometry, we also specify the neural network architecture $(D,L,h)$ and the sizes of the sample and core spaces (|$\absolutevalue{\mathcal{S}}$ | and |$\absolutevalue{\mathcal{V}}$ |) used for that particular calculation. These values, in conjunction with the general methodology and default parameters outlined in Appendix A, define the complete setup for each experiment.
Atom x y z
N2$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
N -0.556 0.0 0.0
N 0.556 0.0 0.0
CH4$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
C 0.0 0.0 0.0
H 0.5541 0.7996 0.4965
H 0.6833 -0.8134 -0.2536
H -0.7782 -0.3735 0.6692
H -0.4593 0.3874 -0.9121
LiF$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
F 2.0 0.0 0.0
Li 3.0 0.0 0.0
CH2O$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
O 0.6123 0.0 0.0
C -0.6123 0.0 0.0
H -1.2 0.2426 -0.8998
H -1.2 -0.2424 0.8998
LiCl$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
Cl 2.0 0.0 0.0
Li 3.0 0.0 0.0
CH4O$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
O 0.7079 0.0 0.0
C -0.7079 0.0 0.0
H -1.0732 -0.769 0.6852
H -1.0731 -0.1947 -1.0113
H -1.0632 0.9786 0.3312
H 0.9936 -0.8804 -0.298
Li2O$(D,L,h)=(1,2,256)$, $|\mathcal{S}|=|\mathcal{V}|=1024$
O 2.866 -0.25 0.0
Li 3.732 0.25 0.0
Li 2.0 0.25 0.0
C2H4O$(D,L,h)=(1,2,512)$, $|\mathcal{S}|=|\mathcal{V}|=4096$
O -0.0007 0.8141 0.0
C 0.7509 -0.4065 0.0
C -0.7502 -0.4076 0.0
H 1.2625 -0.6786 0.9136
H 1.2625 -0.6787 -0.9136
H -1.2614 -0.6806 -0.9136
H -1.2614 -0.6805 0.9136
C2H4O2$(D,L,h)=(1,2,512)$, $|\mathcal{S}|=32768,|\mathcal{V}|=4096$
O -0.3035 1.289 -0.0002
O -0.98 -0.8878 -0.0002
C 1.3743 -0.3516 -0.0002
C -0.0907 -0.0496 0.0006
H 1.8368 0.057 -0.9021
H 1.84 0.0676 0.8952
H 1.5207 -1.4356 0.0064
H -1.2598 1.5081 -0.0008
Table 5: Geometries (in Angstroms) and corresponding training hyperparameters for all molecules studied in Section III.1.1.